OpenStudy (anonymous):

Soybean meal is 18% protein, cornmeal is 9% protein. How many of each should be mixed together in order to get 360lbs mixture that is 14% protein?

OpenStudy (anonymous):

we need the density of the soybean/cornmeal for the 360lbs

OpenStudy (anonymous):

it does not come with density

OpenStudy (anonymous):

x+y=360

OpenStudy (anonymous):

0.06x+0.18y=?

OpenStudy (anonymous):

for final mixture 0.06x+0.18y=0.14(360) initial condition x+y=360 solving these simultaneous equations would get the answer

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

ok, your telling me to solve an equation I dont know how to solve.

OpenStudy (anonymous):

yeah my bad i just copied from post above x- Soybean meal is 18% protein y- cornmeal is 9% protein. we are trying get mix them to get 360lb mixture which means x+y=360 mixture should have 14% protein that explains other equation

OpenStudy (anonymous):

the left side is the protein weight found from adding the percent soybean and cornmeal, the right side is the protein weight found by taking 14% of the total weight of the mixture

OpenStudy (anonymous):

Ok, this question seems impossible to answer

OpenStudy (anonymous):

soybean 200lb cornmeal 160lb